# Latent Plan Transformer for Trajectory Abstraction: Planning as Latent Space Inference

### v3

@@ 5 Related work
Decision-Making via Sequence Modeling Chen et al. (2021) propose Decision Transformer (DT), pioneering this paradigm shift. Concurrently, Janner et al. (2021) explore beam search upon the learned Transformer for model-based planning and inspired later work that searches over the latent state space (Zhang et al., 2022). Lee et al. (2022) report DT’s capability in multi-task setting. Zheng et al. (2022) explore the online extension of DT. Yamagata et al. (2023) augment the Monte Carlo RTG in DT with a Q function and show improvement in tasks requiring trajectory stitching. Janner et al. (2022) explore diffusion models (Ho et al., 2020) as an alternative generative model family for decision-making. Our model differentiates from all above in data specification and model formulation.
Latent Trajectory Variables in Behavior Cloning Yang et al. (2022); Paster et al. (2022) investigate the DT’s overfitting to environment contingencies and propose latent variable solutions. Our model is closely related to theirs but unique in an EM-style algorithm for MLE. Ajay et al. (2021); Lynch et al. (2020) propose latent variable models to make Markovian policies temporally extended. Their models are more related to VAE (Kingma and Welling, 2014).
Offline Reinforcement Learning Since the offline static datasets only partially cover the state transition spaces, efforts from a conventional RL perspective focus on imposing pessimistic biases to value iteration (Kumar et al., 2020; Kostrikov et al., 2021; Uehara and Sun, 2021; Xie et al., 2021; Cheng et al., 2022). Fujimoto and Gu (2021) show that simply augmenting value-based methods with behavior cloning achieves impressive performance. Emmons et al. (2021) report that supervised learning on return-conditioned policies is competitive to value-based methods in offline RL. Our MLE objective is more related to the supervised learning methods. The latent variable inference further imposes temporal consistency, acting as a replacement of value iteration.
Hierarchical RL Methods like OPAL (Ajay et al., 2021), OPOSM (Freed et al., 2023) address TD-learning’s limitations in long-range credit assignment using a two-stage approach: discovering skills from shorter subsequences to reduce the planning horizon, then applying skill-level CQL or online model-based planning on the reduced horizons. This paper focuses on comparing various methods for long-range credit assignment on the original horizon. Future work includes first discovering skills and then modeling them with a skill-level LPT to further extend the effective horizon.
@@ 6.1 Overview
Our empirical study adopts the convention from offline RL. We first train our model with the offline data and then test it as an agent in the corresponding task. More training details and ablation studies of LPT can be found in Sections A.2 and A.4.
OpenAI Gym-Mujoco The D4RL offline RL dataset (Fu et al., 2020) features densely-rewarded locomotion tasks including Halfcheetah, Hopper, and Walker2D. We test for medium and medium-replay. It also includes Antmaze, a locomotion and goal-reaching task with extremely sparse reward. The agent will only receive a reward of 1 if hitting the target location and 0 otherwise. We use its umaze and umaze-diverse variants.
@@
Connect Four This is a tile-based game, where the agent plays against a stochastic opponent (Paster et al., 2022), receiving at the end of an episode 1 reward for winning, 0 for a draw, and -1 for losing.
Baselines We compare the performance of LPT with several representative baselines including CQL  (Kumar et al., 2020), DT  (Chen et al., 2021) and QDT  (Yamagata et al., 2023). CQL baseline results are obtained from Kumar et al. (2020). QDT baseline results are from Yamagata et al. (2023). The DT results for Gym-Mujoco and Maze2D tasks are from Yamagata et al. (2023), Antmaze from Zheng et al. (2022), and Kitchen implemented based on the published source code. CQL and DT results in the Connect Four experiments are from Paster et al. (2022). The mean and standard deviation of our model, shown as LPT and LPT-EI, are reported over 5 seeds.
@@
kitchen-mixed $51.0$ $22.3$ $17.2$ $61.9\pm 1.22$ $\bm{64.7}\pm 0.51$
kitchen-partial $49.8$ $20.4$ $10.5$ $61.2\pm 1.75$ $\bm{65.3}\pm 0.62$
@@ 6.2 Credit assignment
The Gym-Mujoco environment was a standard testbed for high-dimensional continuous control during the development of modern RL algorithms (Lillicrap et al., 2015). In this environment, step-wise rewards were believed to be critical for TD learning methods. In the setup of offline RL, Chen et al. (2021) reported the failure of the competitive CQL baseline when delaying step-wise rewards until the end of the trajectories. DT and QDT are reported to be robust to this alternation. As shown in Table 1, the proposed model, LPT, outperforms these baselines when the data specifications are the same. Notably, LPT even excels in most of the control tasks when compared with the baselines with step-wise rewards.
@@
Maze navigation tasks with fully delayed rewards align with our intuition of a planning problem, for it involves decision-making at certain critical states absent of instantaneous feedback. An ideal planner would take in the expected total return and calculate the sequential decisions, automatically distributing credits from the extremely sparse and fully delayed rewards. According to Yamagata et al. (2023), DT fails in these tasks. Our proposed model LPT outperforms QDT by a large margin in all three variants of the maze task. These results validate our hypothesis that the additional plan prediction KL imposes temporal consistency on autoregressive policies.
@@
Antmaze-umaze 74.074.0{74.0}74.0 $53.3\pm 5.52$ $80.8\pm 4.83$ $\bm{92.4}\pm 0.80$
Antmaze-umaze-diverse 84.084.0{84.0}84.0 $52.5\pm 5.89$ $78.5\pm 1.66$ $\bm{84.4}\pm 1.96$
@@ 6.3 Trajectory stitching
In addition to credit assignment, the setup of offline RL further presents a challenge, trajectory stitching (Fu et al., 2020), which articulates the problem of shifting the trajectory distribution towards sparsely covered regimes with higher returns. In the Franka Kitchen environment, both the mixed, and partial datasets contain undirected data where the robot executes subtasks that do not necessarily achieve the goal configuration. The "mixed" dataset contains no complete solution trajectories, necessitating that the agent learn to piece together relevant sub-trajectories. A similar setting happens in Maze2D domain. Taking Maze2D-medium as an example, in the training set, the average return of all trajectories is $3.98$ with a standard deviation of $10.44$, where the max return is $47$. DT’s score is only marginally above the average return. Yamagata et al. (2023) attribute DT’s failure in Maze2D to its difficulty with trajectory stitching.
@@
Fig. 2 visualizes samples from the training data and successful trajectories in testing. The left panels show that trajectories in training are suboptimal in terms of (1) being short in length and (2) containing very few goal-reaching instances. Trajectories on the right are generated by $10$ random runs with LPT, where the agent successfully navigates to the end goal from random starting positions in an effective manner. This indicates that the agent can discover the correlation between different $y$s to facilitate such stitching.
To probe into the agent’s understanding of trajectories’ returns, we visualize the representation space of the latent variables. The left of Fig. 3 is the aggregated posterior distribution of $z_{0}$. We can see that $z_{0}$ infered from $p_{\theta}(z_{0}|y)$ are distant away from the training population. The agent understands they are not very likely in the training set. The right of Fig. 3 is the distribution of $z$, which is transformed from $z_{0}$ with the UNet, $z=U_{\alpha}(z_{0})$. We observe that $z$s from the generated trajectories become “in-distribution” in the sense that some of them are mingled into the training population and the remaining lie inside a region coverable through linear interpolation of training samples. The agent understands what trajectories to generate even if they are unlikely among what it has seen.
@@ 6.4 Environment contingencies
To live in a stochastic world, contingent planning that is adaptable to unforeseen noises is desirable. Paster et al. (2022); Yang et al. (2022) discover that DT’s performance would degrade in stochastic environments due to inevitable overfitting towards contingencies. We examine LPT and other baselines in Connect Four from Paster et al. (2022). Connect Four is a two-player game, where the opponent will make adversarial moves to deliberately disturb an agent’s plan. According to the empirical study from Paster et al. (2022), the degradation of DT is more significant than in stochastic Gym tasks from Yang et al. (2022). As shown in Table 4, LPT achieves the highest score with minimal variance. The ESPER baseline is from Paster et al. (2022), which is very relevant to LPT as it is also a latent variable model. ESPER learns the latent variable model with an adversarial loss. It further adds a clustering loss in the latent space. LPT’s on-par performance may justify that MLE upon a more flexible prior can play an equal role.
@@ 7 Limitation
We omit the Antmaze-large experiment from the main text and included potential reasons for LPT’s unsatisfactory performance in Section A.3. Another interesting direction is to study LPT’s continual learning potential. During planning, LPT explores with provably efficient posterior sampling (Osband et al., 2013; Osband and Van Roy, 2017).
@@ 8 Summary
We study generative modeling for planning in the absence of step-wise rewards. We propose LPT which generates trajectory and return from a latent variable. In learning, posterior sampling of the latent variable naturally gathers sub-trajectories to form an episode-wise abstraction despite finite context in training. In inference, the posterior sampling given the target final return explores the optimal regime of the latent space. It produces a latent variable that guides the autoregressive policy to execute consistently. Across diverse evaluations, LPT demonstrates competitive capacities of nuanced credit assignments, trajectory stitching, and adaptation to environmental contingencies. Contemporary work extends LPT’s application to online molecule design (Kong et al., 2024). Future research directions include studying online and multi-agent variants of this model, exploring its application in real-world robotics, and investigating its potential in embodied agents.
@@ A.1 Details about model and learning
Given a trajectory $\tau$, $z\in\mathbb{R}^{d}$ is the latent vector to represent the variable-length trajectory. $y\in\mathbb{R}$ is the return of the trajectory. With offline training trajectory-return pairs $\{(\tau_{i},y_{i}),i=1,...,n\}$. The log-likelihood function is $L(\theta)=\sum_{i=1}^{n}\log p_{\theta}(\tau_{i},y_{i})$, with learning gradient $\nabla_{\theta}L(\theta)=\sum_{i=1}^{n}\nabla_{\theta}\log p_{\theta}(\tau_{i},y_{i})$. We derive the form of $\nabla_{\theta}\log p_{\theta}(\tau_{i},y_{i})$, proving Eq. 4 below, dropping index subscript i for simplicity.
@@
[…]
@@ A.3 Discussion on data quality of Antmaze medium and large
Fig. 4 plots the distributions of final returns and the trajectory lengths. Surprisingly, this dataset consists of 5448 trajectories (75.86%) with length=1, 893 trajectories (12.43%) with length=1000, and only 841 trajectories (11.71%) with lengths in between. Such a biased trajectory coverage can be detrimental to sequence models like LPT, which learn to make decisions by discovering correlations between trajectories and returns.
As a reference, Fig. 5 shows the distributions of final returns and the trajectory lengths of Maze2D-large, a task where LPT performs well. It is important to note that TD-learning methods, such as CQL and QDT, rely solely on $(s,a,s^{\prime},r)$ tuples and are less affected by the trajectory length distribution in the dataset. Consequently, Antmaze-large in D4RL remains a fair dataset for these methods to perform offline RL.
@@ A.4 Ablation study
A.5 Continual learning with online data
We are also interested in LPT’s potential in finetuning or even continual learning. Inspired by ODT (Zheng et al., 2022), we employ a trajectory replay butter (Mnih et al., 2015) to store samples from online interaction in a first-in-first-out manner. After the completion of each episode, we update LPT with the same learning algorithm as with the offline data. Note that ODT introduces some techniques additional to DT. In contrast, LPT explores with the provably efficient posterior sampling (Osband et al., 2013; Osband and Van Roy, 2017). We report the results in Table 11. Despite the significance in a few tasks, the improvement is within 1 standard deviation of the mean for the majority. We observe a similar pattern in ODT (Zheng et al., 2022).
Table 11: Evaluation results of online Open AI Gym MuJoCo and Antmaze tasks. ODT baselines are sourced from Zheng et al. (2022). Our results are reported over $5$ seeds.
Dataset Step-wise Reward Final Return
ODT $\Delta$ LPT $\Delta$
halfcheetah-medium $42.16\pm 1.48$ $-0.56$ $43.26\pm 0.59$ $0.13$
halfcheetah-medium-replay $40.2\pm 1.61$ $0.43$ $40.63\pm 0.28$ $0.99$
hopper-medium $97.54\pm 2.1$ $30.59$ $64.84\pm 10.29$ $6.32$
hopper-medium-replay $88.89\pm 6.33$ $2.25$ $72.44\pm 8.07$ $1.27$
walker2d-medium $76.79\pm 2.30$ $4.60$ $79.54\pm 5.11$ $1.69$
walker2d-medium-replay $76.86\pm 4.04$ $7.94$ $78.99\pm 3.84$ $6.68$
Antmaze-umaze $88.5\pm 5.88$ $35.4$ $83.5\pm 3.28$ $22.3$
Antmaze-diverse $56.00\pm 5.69$ $5.8$ $75.6\pm 2.8$ $8.0$

### v4

@@ 5 Related work
Decision-Making via Sequence Modeling Chen et al. [2021] propose Decision Transformer (DT), pioneering this paradigm shift. Concurrently, Janner et al. [2021] explore beam search upon the learned Transformer for model-based planning and inspired later work that searches over the latent state space [Zhang et al., 2022]. Lee et al. [2022] report DT’s capability in multi-task setting. Zheng et al. [2022] explore the online extension of DT. Yamagata et al. [2023] augment the Monte Carlo RTG in DT with a Q function and show improvement in tasks requiring trajectory stitching. Janner et al. [2022] explore diffusion models [Ho et al., 2020] as an alternative generative model family for decision-making. Our model differentiates from all above in data specification and model formulation.
Latent Trajectory Variables in Behavior Cloning Yang et al. [2022], Paster et al. [2022] investigate the DT’s overfitting to environment contingencies and propose latent variable solutions. Our model is closely related to theirs but unique in an EM-style algorithm for MLE. Ajay et al. [2021], Lynch et al. [2020] propose latent variable models to make Markovian policies temporally extended. Their models are more related to VAE [Kingma and Welling, 2014].
Offline Reinforcement Learning Since the offline static datasets only partially cover the state transition spaces, efforts from a conventional RL perspective focus on imposing pessimistic biases to value iteration [Kumar et al., 2020, Kostrikov et al., 2021, Uehara and Sun, 2021, Xie et al., 2021, Cheng et al., 2022]. Fujimoto and Gu [2021] show that simply augmenting value-based methods with behavior cloning achieves impressive performance. Emmons et al. [2021] report that supervised learning on return-conditioned policies is competitive to value-based methods in offline RL. Our MLE objective is more related to the supervised learning methods. The latent variable inference further imposes temporal consistency, acting as a replacement of value iteration.
Hierarchical RL Methods like OPAL [Ajay et al., 2021], OPOSM [Freed et al., 2023] address TD-learning’s limitations in long-range credit assignment using a two-stage approach: discovering skills from shorter subsequences to reduce the planning horizon, then applying skill-level CQL or online model-based planning on the reduced horizons. This paper focuses on comparing various methods for long-range credit assignment on the original horizon. Future work includes first discovering skills and then modeling them with a skill-level LPT to further extend the effective horizon.
@@ 6.1 Overview
Our empirical study adopts the convention from offline RL. We first train our model with the offline data and then test it as an agent in the corresponding task. More training details and ablation studies of LPT can be found in Sections˜A.2 and A.4.
OpenAI Gym-Mujoco The D4RL offline RL dataset [Fu et al., 2020] features densely-rewarded locomotion tasks including Halfcheetah, Hopper, and Walker2D. We test for medium and medium-replay. It also includes Antmaze, a locomotion and goal-reaching task with extremely sparse reward. The agent will only receive a reward of 1 if hitting the target location and 0 otherwise. We use its umaze and umaze-diverse variants.
@@
Connect Four This is a tile-based game, where the agent plays against a stochastic opponent [Paster et al., 2022], receiving at the end of an episode 1 reward for winning, 0 for a draw, and -1 for losing.
Baselines We compare the performance of LPT with several representative baselines including CQL  [Kumar et al., 2020], DT  [Chen et al., 2021] and QDT  [Yamagata et al., 2023]. CQL baseline results are obtained from Kumar et al. [2020]. QDT baseline results are from Yamagata et al. [2023]. The DT results for Gym-Mujoco and Maze2D tasks are from Yamagata et al. [2023], Antmaze from Zheng et al. [2022], and Kitchen implemented based on the published source code. CQL and DT results in the Connect Four experiments are from Paster et al. [2022]. The mean and standard deviation of our model, shown as LPT and LPT-EI, are reported over 5 seeds.
@@
kitchen-mixed $51.0$ $22.3$ - - $17.2$ - $61.9\pm 1.22$ $\bm{64.7}\pm 0.51$
kitchen-partial $49.8$ $20.4$ - - $10.5$ - $61.2\pm 1.75$ $\bm{65.3}\pm 0.62$
@@ 6.2 Credit assignment
The Gym-Mujoco environment was a standard testbed for high-dimensional continuous control during the development of modern RL algorithms [Lillicrap et al., 2015]. In this environment, step-wise rewards were believed to be critical for TD learning methods. In the setup of offline RL, Chen et al. [2021] reported the failure of the competitive CQL baseline when delaying step-wise rewards until the end of the trajectories. DT and QDT are reported to be robust to this alternation. As shown in Table 1, the proposed model, LPT, outperforms these baselines when the data specifications are the same. Notably, LPT even excels in most of the control tasks when compared with the baselines with step-wise rewards.
@@
Maze navigation tasks with fully delayed rewards align with our intuition of a planning problem, for it involves decision-making at certain critical states absent of instantaneous feedback. An ideal planner would take in the expected total return and calculate the sequential decisions, automatically distributing credits from the extremely sparse and fully delayed rewards. According to Yamagata et al. [2023], DT fails in these tasks. Our proposed model LPT outperforms QDT by a large margin in all three variants of the maze task. These results validate our hypothesis that the additional plan prediction KL imposes temporal consistency on autoregressive policies.
@@
Antmaze-umaze 74.0{74.0} $53.3\pm 5.52$ $80.8\pm 4.83$ $\bm{92.4}\pm 0.80$
Antmaze-umaze-diverse 84.0{84.0} $52.5\pm 5.89$ $78.5\pm 1.66$ $\bm{84.4}\pm 1.96$
@@ 6.3 Trajectory stitching
In addition to credit assignment, the setup of offline RL further presents a challenge, trajectory stitching [Fu et al., 2020], which articulates the problem of shifting the trajectory distribution towards sparsely covered regimes with higher returns. In the Franka Kitchen environment, both the mixed, and partial datasets contain undirected data where the robot executes subtasks that do not necessarily achieve the goal configuration. The "mixed" dataset contains no complete solution trajectories, necessitating that the agent learn to piece together relevant sub-trajectories. A similar setting happens in Maze2D domain. Taking Maze2D-medium as an example, in the training set, the average return of all trajectories is $3.98$ with a standard deviation of $10.44$, where the max return is $47$. DT’s score is only marginally above the average return. Yamagata et al. [2023] attribute DT’s failure in Maze2D to its difficulty with trajectory stitching.
@@
Fig.˜2 visualizes samples from the training data and successful trajectories in testing. The left panels show that trajectories in training are suboptimal in terms of (1) being short in length and (2) containing very few goal-reaching instances. Trajectories on the right are generated by $10$ random runs with LPT, where the agent successfully navigates to the end goal from random starting positions in an effective manner. This indicates that the agent can discover the correlation between different $y$s to facilitate such stitching.
To probe into the agent’s understanding of trajectories’ returns, we visualize the representation space of the latent variables. The left of Fig.˜3 is the aggregated posterior distribution of $z_{0}$. We can see that $z_{0}$ infered from $p_{\theta}(z_{0}|y)$ are distant away from the training population. The agent understands they are not very likely in the training set. The right of Fig.˜3 is the distribution of $z$, which is transformed from $z_{0}$ with the UNet, $z=U_{\alpha}(z_{0})$. We observe that $z$s from the generated trajectories become “in-distribution” in the sense that some of them are mingled into the training population and the remaining lie inside a region coverable through linear interpolation of training samples. The agent understands what trajectories to generate even if they are unlikely among what it has seen.
@@ 6.4 Environment contingencies
To live in a stochastic world, contingent planning that is adaptable to unforeseen noises is desirable. Paster et al. [2022], Yang et al. [2022] discover that DT’s performance would degrade in stochastic environments due to inevitable overfitting towards contingencies. We examine LPT and other baselines in Connect Four from Paster et al. [2022]. Connect Four is a two-player game, where the opponent will make adversarial moves to deliberately disturb an agent’s plan. According to the empirical study from Paster et al. [2022], the degradation of DT is more significant than in stochastic Gym tasks from Yang et al. [2022]. As shown in Table 4, LPT achieves the highest score with minimal variance. The ESPER baseline is from Paster et al. [2022], which is very relevant to LPT as it is also a latent variable model. ESPER learns the latent variable model with an adversarial loss. It further adds a clustering loss in the latent space. LPT’s on-par performance may justify that MLE upon a more flexible prior can play an equal role.
@@ 7 Limitation
We omit the Antmaze-large experiment from the main text and included potential reasons for LPT’s unsatisfactory performance in Section˜A.3. Another interesting direction is to study LPT’s continual learning potential. During planning, LPT explores with provably efficient posterior sampling [Osband et al., 2013, Osband and Van Roy, 2017].
@@ 8 Summary
We study generative modeling for planning in the absence of step-wise rewards. We propose LPT which generates trajectory and return from a latent variable. In learning, posterior sampling of the latent variable naturally gathers sub-trajectories to form an episode-wise abstraction despite finite context in training. In inference, the posterior sampling given the target final return explores the optimal regime of the latent space. It produces a latent variable that guides the autoregressive policy to execute consistently. Across diverse evaluations, LPT demonstrates competitive capacities of nuanced credit assignments, trajectory stitching, and adaptation to environmental contingencies. Contemporary work extends LPT’s application to online molecule design [Kong et al., 2024]. Future research directions include studying online and multi-agent variants of this model, exploring its application in real-world robotics, and investigating its potential in embodied agents.
@@ A.1 Details about model and learning
Given a trajectory $\tau$, $z\in\mathbb{R}^{d}$ is the latent vector to represent the variable-length trajectory. $y\in\mathbb{R}$ is the return of the trajectory. With offline training trajectory-return pairs $\{(\tau_{i},y_{i}),i=1,...,n\}$. The log-likelihood function is $L(\theta)=\sum_{i=1}^{n}\log p_{\theta}(\tau_{i},y_{i})$, with learning gradient $\nabla_{\theta}L(\theta)=\sum_{i=1}^{n}\nabla_{\theta}\log p_{\theta}(\tau_{i},y_{i})$. We derive the form of $\nabla_{\theta}\log p_{\theta}(\tau_{i},y_{i})$, proving Eq.˜4 below, dropping index subscript i for simplicity.
@@
$=\frac{1}{p_{\theta}(\tau,y)}\int\nabla_{\theta}p_{\theta}(\tau,y,z=U_{\alpha}(z_{0}))dz_{0}$
$=\int\frac{p_{\theta}(\tau,y,z=U_{\alpha}(z_{0}))}{p_{\theta}(\tau,y)}\nabla_{\theta}\log p_{\theta}(\tau,y,z=U_{\alpha}(z_{0}))dz_{0}$
$=\int p_{\theta}(z_{0}|\tau,y)\nabla_{\theta}\log p_{\theta}(\tau,y,z=U_{\alpha}(z_{0}))dz_{0}$
@@ A.3 Discussion on data quality of Antmaze medium and large
Fig.˜4 plots the distributions of final returns and the trajectory lengths. Surprisingly, this dataset consists of 5448 trajectories (75.86%) with length=1, 893 trajectories (12.43%) with length=1000, and only 841 trajectories (11.71%) with lengths in between. Such a biased trajectory coverage can be detrimental to sequence models like LPT, which learn to make decisions by discovering correlations between trajectories and returns.
As a reference, Fig.˜5 shows the distributions of final returns and the trajectory lengths of Maze2D-large, a task where LPT performs well. It is important to note that TD-learning methods, such as CQL and QDT, rely solely on $(s,a,s^{\prime},r)$ tuples and are less affected by the trajectory length distribution in the dataset. Consequently, Antmaze-large in D4RL remains a fair dataset for these methods to perform offline RL.